### v3

@@ 7.1 Challenges
Insufficient Mathematical Capability. LLMs struggle with tasks requiring precise calculations or logical reasoning, such as multi-robot path planning or trajectory optimization. This limitation reduces their effectiveness in scenarios where quantitative accuracy is critical. Mirzadeh et al. [67] performed a detailed comparison and investigation on the mathematical understanding and problem-solving ability of several state-of-art LLMs. Specifically, LLMs exhibit noticeable variance when responding to different variations of the same question, with performance declining significantly when only the numerical values are altered. Furthermore, their reasoning capabilities are fragile; they often mimic patterns observed in training data rather than performing genuine logical deduction. This fragility is exacerbated by an increase in the number of clauses within a question, even when the added clauses are irrelevant to the reasoning chain, leading to performance drops of up to 65% in state-of-the-art models. These vulnerabilities present serious challenges for MRS, where precise calculations and robust reasoning are essential for collision-free trajectories, spatial planning, and efficient task execution. Addressing these limitations is critical for deploying LLMs reliably in mathematically intensive applications.
@@
Difficulties in Field Deployment. Current options for using LLMs include server-based models, which are usually close-sourced, and open-source models that individuals can deploy locally. Examples of server-based models include OpenAI GPT [1], Anthropic Claude [7], and Google Gemini (formerly known as Bard) [27], and open-sourced LLMs that individuals can run locally includes Meta Llama [19], Falcon [3], Alibaba Qwen [101], and DeepSeek V3 [53] and R1 [29]. The server-based models require a reliable internet connection to send inquiries and receive responses, thus making deploying the MRS with LLMs in remote locations unachievable, which is typical for field robot systems. Moreover, server-based LLMs heavily rely on the performance of the server, where the server outage can interrupt the systems built on LLMs entirely. This issue is especially vital for multi-robot teams as the LLM guides the inter-robot collaboration and decision-making. On the other hand, local models can avoid using the servers but require hardware onboard that is powerful enough to run LLMs locally.
@@ 7.2 Opportunities
Advanced Reasoning Techniques. Improving the reasoning capabilities of LLMs is critical for addressing their current limitations in logical and mathematical tasks. Techniques such as Chain of Thought (CoT) prompting, fine-tuning with explicit reasoning labels, integrating symbolic reasoning, and training with RL can enhance the ability of LLMs to handle complex multi-step problems. By advancing reasoning methods, LLMs can better support tasks that require precision and logical deduction, such as multi-robot path planning and coordination.
Task-specific and Lightweight Models. While large-scale LLMs offer superior performance, they are often impractical for resource-constrained environments. Developing task-specific and lightweight models tailored for multi-robot applications can mitigate this issue. Models like SmolVLM, Moondream 2B, PaliGemma 3B, and Qwen2-VL 2B demonstrate how smaller architectures can reduce computational demands and latency while maintaining adequate performance for specific tasks. Model distillation is another approach to make small models more capable by distilling knowledge from a more capable LLM, like DeepSeek-R1-Distill-Qwen-1.5B, where the knowledge from DeepSeek R1 is distilled into a small Qwen2.5-Math-1.5B model. Balancing efficiency and effectiveness is key to enabling scalable deployments of LLMs in field robotics.
@@
Latest More Capable LLMs. The continued development of state-of-the-art LLMs opens new possibilities for MRS. Models such as PaliGemma, Qwen, GPT o3 (mini), and DeepSeek V3 and R1 offer enhanced reasoning, understanding, and multitasking capabilities. Incorporating these advanced models into MRS research can accelerate progress by providing improved baseline performance and enabling innovative applications. Exploring their integration with robotics systems can further push the boundaries of what multi-robot teams can achieve.
@@ 9 Acknowledgment
10 Competing interests
The authors declare that they have no competing interests.

### v4

@@ 7.1 Challenges
Insufficient Mathematical Capability. LLMs struggle with tasks requiring precise calculations or logical reasoning, such as multi-robot path planning or trajectory optimization. This limitation reduces their effectiveness in scenarios where quantitative accuracy is critical. Mirzadeh et al. [87] performed a detailed comparison and investigation on the mathematical understanding and problem-solving ability of several state-of-the-art LLMs. Specifically, LLMs exhibit noticeable variance when responding to different variations of the same question, with performance declining significantly when only the numerical values are altered. Furthermore, their reasoning capabilities are fragile; they often mimic patterns observed in training data rather than performing genuine logical deduction. This fragility is exacerbated by an increase in the number of clauses within a question, even when the added clauses are irrelevant to the reasoning chain, leading to performance drops of up to 65% in state-of-the-art models. These vulnerabilities present serious challenges for MRS, where precise calculations and robust reasoning are essential for collision-free trajectories, spatial planning, and efficient task execution. Addressing these limitations is critical for deploying LLMs reliably in mathematically intensive applications.
@@
Difficulties in Field Deployment. Current options for using LLMs include server-based models, which are usually closed-source, and open-source models that individuals can deploy locally. Examples of server-based models include OpenAI GPT [1], Anthropic Claude [7], and Google Gemini (formerly known as Bard) [32], and open-sourced LLMs that individuals can run locally includes Meta Llama [23], Falcon [3], Alibaba Qwen [131], and DeepSeek V3 [70] and R1 [34]. The server-based models require a reliable internet connection to send inquiries and receive responses, thus making deploying the MRS with LLMs in remote locations unachievable, which is typical for field robot systems. Moreover, server-based LLMs heavily rely on the performance of the server, where a server outage can interrupt the systems built on LLMs entirely. This issue is especially vital for multi-robot teams as the LLM guides the inter-robot collaboration and decision-making. On the other hand, local models can avoid using the servers but require hardware onboard that is powerful enough to run LLMs locally.
@@
Multi-modal Integration Difficulties. While VLMs and VLAs extend the role of LLMs in MRS by grounding reasoning in perception and linking directly to action, they also introduce new challenges. Multi-modal inference often incurs higher latency due to the added cost of processing images, video, or sensor streams alongside text [65]. Domain shift remains a significant issue, as models trained on internet-scale vision-language corpora may not generalize reliably to the specialized sensing modalities of robots (e.g., LiDAR, thermal cameras, aerial views) [82]. Moreover, fusing heterogeneous and asynchronous sensor data across multiple robots introduces synchronization and consistency problems, especially in decentralized settings [82]. VLAs face the additional challenge of grounding action outputs across diverse robot morphologies, making it difficult to generalize learned behaviors across heterogeneous teams. Addressing these multi-modal integration difficulties is critical for ensuring that perception-grounded LLMs can operate robustly in real-world MRS deployments.
@@ 7.2 Opportunities
Advanced Reasoning Techniques. Improving the reasoning capabilities of LLMs is critical for addressing their current limitations in logical and mathematical tasks. Techniques such as Chain of Thought (CoT) prompting [122], fine-tuning with explicit reasoning labels [138], integrating symbolic reasoning, and training with RL can enhance the ability of LLMs to handle complex multi-step problems. By advancing reasoning methods, LLMs can better support tasks that require precision and logical deduction, such as multi-robot path planning and coordination.
Task-specific and Lightweight Models. While large-scale LLMs offer superior performance, they are often impractical for resource-constrained environments. Developing task-specific and lightweight models tailored for multi-robot applications can mitigate this issue [9]. Models like SmolVLM, Moondream 2B, PaliGemma-2 3B, and Qwen2-VL 2B demonstrate how smaller architectures can reduce computational demands and latency while maintaining adequate performance for specific tasks. Model distillation is another approach to make small models more capable by distilling knowledge from a more capable LLM, like DeepSeek-R1-Distill-Qwen-1.5B, where the knowledge from DeepSeek R1 is distilled into a small Qwen2.5-Math-1.5B model [128]. Balancing efficiency and effectiveness is key to enabling scalable deployments of LLMs in field robotics.
Multi-modal and Embodied Extensions. Emerging research on multi-modal LLMs suggests promising opportunities for extending MRS capabilities. Video-language transformers, point cloud-language models [26], and spatio-temporal VLAs could enable richer reasoning over dynamic events [140], 3D spatial layouts, and real-time cooperative interactions. These architectures would allow multi-robot teams to better interpret evolving environments and adapt their strategies accordingly. As multi-modal and embodied extensions mature, they could bridge the gap between high-level symbolic reasoning and low-level perception-action execution, unlocking more adaptive and autonomous multi-robot coordination [9].
@@
Latest More Capable LLMs. The continued development of state-of-the-art LLMs opens new possibilities for MRS. Models such as PaliGemma-2, Qwen-3, Claude 4 Sonnet/Opus, GPT o3 or o4-mini, GPT-5, and DeepSeek V3 and R1 offer enhanced reasoning, understanding, and multitasking capabilities [86, 126]. Incorporating these advanced models into MRS research can accelerate progress by providing improved baseline performance and enabling innovative applications. Exploring their integration with robotics systems can further push the boundaries of what multi-robot teams can achieve.